Use inductive reasoning to predict the most probable next number in te following lists. 2, 5, 10, 17, 26, (?)



1
Expert's answer
2021-09-21T12:30:39-0400

"a_2-a_1=5-2=3"

"a_3-a_2=10-5=5"

"a_4-a_3=17-10=7"

"a_5-a_4=26-17=9"

So, the most probable "a_6=a_5+11=26+11=37."